OpenStudy (anonymous):

Find the domain and range of the real function. f(x) = √x-1.

Parth (parthkohli):

If the function is real, then you can't have an imaginary square root. A square root is imaginary when the thing inside it is negative.

Parth (parthkohli):

So, you only start with 0. You can't have things such as\[\sqrt{-1}, \sqrt{-2}\cdots \]

Parth (parthkohli):

So,\[ x - 1\ge0\]Solve that inequality. When you do, we'd sort the domain and range out too.

Parth (parthkohli):

@ashna You there?

OpenStudy (anonymous):

x^2 - 1 = 0

OpenStudy (anonymous):

x - 1 > 0

Parth (parthkohli):

What? Just solve the equality \(x - 1\ge 0\)

OpenStudy (anonymous):

x > 1

Parth (parthkohli):

inequality*

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

what i did is wrong ?

Parth (parthkohli):

So, \(x\) needs to be greater than or equal to 1. That's just natural numbers. Can be written as:\[\text{Domain:} x\in \mathbb{N} \]

Parth (parthkohli):

No, it is correct.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

and range is also greater than or equal to zero right ?

Parth (parthkohli):

Wait a second. Is the question like this?\[ f(x) = \sqrt{x} - 1\]

Parth (parthkohli):

I was correct the first time. Never mind the second solution.

OpenStudy (anonymous):

okay :)

Parth (parthkohli):

Yes, @ashna You are correct for the range too!

OpenStudy (anonymous):

oh okay then .. tnx again :)

Parth (parthkohli):

Can you please confirm if it is \(\sqrt{x}-1\) or \(\sqrt{x - 1}\)?

Parth (parthkohli):

I have given the solution for the latter.

OpenStudy (anonymous):

whole root x - 1